Lakeside Components wishes to purchase parts in one month for sale in the next. On June 1, the company has 10,000 parts in stock, although sales for June are estimated to total 11,200 parts. Total sales of parts are expected to be 12,200 in July and 10,400 in August.

Parts are purchased at a wholesale price of $10. The supplier has a financing arrangement by which Lakeside Components pays 60 percent of the purchase price in the month when the parts are delivered and 40 percent in the following month. Lakeside purchased 15,000 parts in May.


Required:
a.
Estimate purchases (in units) for June and July.
b. Estimate the cash required to make purchases in June and July.
Answer #1

a.

Purchases = Sales - ending units - beginning units

June = 11,200 + 12,200 - 10,000 =13,400

July = 12,200 + 10,400 - 12,200 = 10,400

b.

June payment = (15,000 X $10 X 40%) + (13,400 X $10 X 60%) = $140,400

July payment = (13,400 X $10 X 40%) + (10,400 X $10 X 60%) = $116,000